# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v3

@@ 5.2 Task Sampling Strategy Design
In robust active task sampling, existing strategies evaluate task batches to rank their difficulties in adaptation and then prioritize challenging subsets for optimization 20, 16. Besides the expensive evaluation cost, these strategies are weak in the efficient exploration of the task space.
As Theorem 1 has established the theoretical foundation of amortized evaluation, this necessitates the development of the risk learner from cumulated risk histories. With the model predictive results as amortized evaluation, specific rules can be incorporated into the acquisition function for active sampling.
Meanwhile, it is fascinating for the risk learner to evaluate the machine learner’s adaptation to arbitrarily many tasks with minimal computational cost. Hence, we can easily enlarge the pseudo batch size $\hat{\mathcal{B}}$ for more selection candidates and exploit the epistemic uncertainty from the risk learner, encouraging more exploration in the task space.
@@
After obtaining $p_{\bm{\psi}}(\ell|\bm{\tau},H_{1:t})$, we draw up a batch sampling strategy on the basis of its quantified statistics. The criteria resembles the acquisition function in classical Bayesian optimization (BO), which includes a collection of available evaluation principles, such as expected improvement 57, output information theoretical index 58 or UCB 59.
However, it is also necessary to clarify that the search space is on the sequential task batch instead of machine learners’ parameters, which differs from the ultimate purpose in BO. Central to our approach is the principle of optimism in the face of uncertainty (OFU) 60. We consider the difficult task’s prioritization for robustness and the epistemic uncertainty as pivotal elements in developing acquisition functions. The grounds behind this idea are that (i) the subset with the worst performance deserves extra attention in optimization for adaptation robustness, and (ii) task regions with high predictive uncertainty tend to be underexplored in the last few iterations.
As a result, we present the acquisition function built on the UCB principle 39:
@@
Then, the Simulate-Rank-Filter operation in Eq. (8)c arrives at the task batch for $(t+1)$-th iteration, i.e., $\mathcal{T}_{t+1}^{\mathcal{B}}=\arg\max_{\mathcal{T}^{\mathcal{B}}\subseteq\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}:|\mathcal{T}^{\mathcal{B}}|=\mathcal{B}}\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$. This characterizes the step of the active subset selection from $\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}$, the randomly sampled identifier candidate set with $|\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}|=\hat{\mathcal{B}}$.
In an implementation, we still perform random sampling from $p(\tau)$ and forecast the task-wise acquisition score $a(\cdot)$ from the risk learner. Candidates in $-\mathcal{B}$ acquisition scores are screened to formulate the task batch $\mathcal{T}_{t+1}^{\mathcal{B}}$ for episodic optimization, as illustrated in Fig. 1d. These steps approximately solve Eq. (8)c and obtain $\mathcal{T}_{t+1}^{\mathcal{B}}$ in a heuristic way.
@@ 5.3 Sequentially Optimize the Adaptive Machine Learner
Given the screened $\mathcal{T}_{t+1}$, we execute optimization to update the machine learner’s parameters. The task-specific adaptation risk in $(t+1)$-th iteration is written as $\ell_{t+1,i}(\bm{\theta})$ for the selected task $\bm{\tau}_{i}$. The developed MPTS is agnostic to any-shot learning methods, and the following includes the standard update rule for zero-shot, few-shot, and many-shot scenarios.
@@
The zero-shot setup does not require the support dataset to identify the task. Hence, taking the vanilla DR 43 as an instantiation, we can obtain the update rule as:
@@
Still, we take the typical optimization-based method MAML 28 as an instantiation and provide the update rule as follows:
@@
Machine Learner Updates in Many-Shot Adaptation:
Here, we take finetuning pretrained models to downstream tasks 61 as an instantiation. In this case, each data point $[\bm{x},\bm{y}]$ can be viewed as a task with either its embedding $\bm{\tau}$ or $\bm{x}$ as the task identifier. Then the model update rule can be:
@@ 5.4 Overall Algorithm and Interpretation
For the sake of generality, we write the general form of MPTS in Algorithm 1.
Implementation pipelines in any-shot adaptation.
Putting all the ingredients and optimization recipes together, we can provide some examples of implementation. Still, we take some typical adaptation methods and describe pipelines in the zero-shot, few-shot, and many-shot scenarios. This results in Algorithm 2-7. Since the first iteration in Algorithm 2/4/6 does not involve active sampling, due to no latest history, and the task batch follows the standard random sampling setup.
@@
Output : Selected task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
@@
       $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
       $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
@@
// Simulating Zero-shot, Few-shot, and Many-shot Adaptation Results
@@
Return the screened task batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
@@
Intuitively, MPTS resembles model predictive control (MPC) 62 when treating task sampling under some criteria as an optimal planning problem. In this case, the episodic learning process specifies an underlying dynamical system for MPTS to predict with only one future time step in the simulation to assess the influence of selecting the task batch, and the feedback as exact adaptation risk information further helps improve the episodic risk prediction system. In addition, through the lens of sequential decision-making, we can interpret the optimization pipeline of MPTS from the actor-critic framework in RL 63. In detail, the risk learner works as the critic that predicts adaptation performance in the task $\tau$ given a fixed machine learner. Accordingly, the actor plays the role of selecting the task batch from the acquisition function and then executing the machine learner’s optimization. These two roles are entangled in the MPTS pipeline to achieve robust yet efficient adaptation.
@@
Supplementary Materials for Model Predictive Task Sampling in Any-Shot Adaptation
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
@@
Algorithm 6 MPTS for Pretrained Model Finetuning (Many-Shot Scenarios)
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
@@ Appendix A Quick Guideline to MPTS
Task episodic learning serves as a cornerstone in developing adaptive models by structuring diverse, context-rich learning experiences. One of the pivotal insights underpinning this process is the neural scaling law, which establishes a relationship between task volume, model complexity, and computational resources, offering a principled insight into training foundation models at a certain budget. Recent viewpoints have also shed light on the importance of task quality 17, 29, 19, 30, 20, 64, 18, 65, prompting innovative data curation strategies to refine datasets for pretraining, meta-training, and post-training. Evidence suggests that carefully curated data can significantly reduce task sampling complexity, decrease computational demands, and enhance robustness against distributional shifts—sometimes achieving these goals simultaneously. Despite these advancements, a practical operation such as Evaluate-Rank-Filter still faces challenges associated with costly evaluations from intensive task queries, computational overhead, and massive annotations. Addressing these bottlenecks remains essential to fully realize the potential of task episodic learning in robust efficient foundation model training.
@@
The involvement of the risk learner inevitably brings extra computational overhead in optimization. However, the risk learner used in this work is lightweight with the model complexity $\mathcal{O}(|\bm{\phi}|+|\bm{\psi}|)<<\mathcal{O}(|\bm{\theta}|)$. We can roughly estimate these extra computations that arise from the predictive model as $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|)T_{\text{MPTS}}\big{)}$ throughout the training phase. Moreover, the computational and task evaluation complexities of different methods are estimated in Table 2. Compared with DRM, MPTS retains more computational and task efficiency when the filtering ratio $\hat{\alpha}$ is high, and the machine learner $\bm{\theta}$ is largely given similar convergence iteration steps. However, exactly evaluating task difficulties from either losses or gradients brings extra computational and annotation overhead.
@@
Among MPTS’s core components, the risk learner works to predict the adaptation risk values based on historical information and further serves the calculation of acquisition functions. Importantly, this work primarily investigates the feasibility and effectiveness of MPTS and does not impose rigid constraints on the form of the risk learner $p(\ell|\bm{\tau},H_{1:t})$ too much in modeling. The design of this risk learner $p(\ell|\bm{\tau},H_{1:t})$ must meet several criteria: it is tractable in optimization, can process historical risk information, and offers uncertainty in prediction.
A series of candidate probabilistic models exist that probably apply to adaptation risk modeling. One alternative choice can be the Gaussian process 66, which provides an analytical form of the predictive distribution. However, its implementation (i) is less scalable in the case of relatively higher dimensional task identifiers, (ii) holds the cubic runtime complexity in obtaining the predictive covariance matrix, (iii) is sensitive to kernel selection, coupled with limited expressiveness of the Gaussian distribution in learned risk functions. Hence, for simplicity and computational efficiency, we adopt the basic VAE-like model and execute a handful of gradient updates to train the risk learner. We leave more advanced risk learner modeling for future exploration.
@@
This work relates to active sampling and Bayesian optimization. The purpose of BO 67 is to sequentially find a global optimum of a black-box function $f(\bm{x})$ expensive to evaluate in $\mathcal{S}$, namely $\bm{x}_{*}=\arg\max_{\bm{x}\in\mathcal{S}\subset\mathbb{R}^{d}}f(\bm{x})$.
In each iteration $t=1,\dots,T$, the BO method actively queries $\bm{x}_{t}$ to evaluate $f(\bm{x}_{t})$, yeilding an output $\ell_{t}=f(\bm{x}_{t})+\epsilon$ with a white noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Due to the high cost of function evaluation, the key to BO is constructing a surrogate model to guide the data point to query. The resulting acquisition function 68 works as an active sampling objective to maximize and obtain the candidate $\bm{x}_{t}$ based on the previous sequence. BO requires limited function evaluations as observations and exploits the correlations in queried data points. These properties make it more theoretically data efficient than random or grid search in seeking the optimal solution 69. This work differs from standard BO as task episodic learning is not the optimal parameter search problem.
@@
The main paper provides the workflow of MPTS in Algorithm 1. For separate scenarios, we attach detailed pseudo algorithms as follows. These illustrated Algorithms are in the context of supervised learning. Regarding RL scenarios, such as meta RL and DR, there is a slight modification for MPTS. As simply picking up worst-case MDPs restricts the task subspace in optimization 33, we adopt the mixture of the identifier subset from the random sampler and the identifier subset from the MPTS sampler. For example, in meta RL, with the pseudo batch size ^$\hat{\mathcal{B}}=1.5\mathcal{B}$, there $1.5\mathcal{B}$ identifier candidates from the random sampler. We retain $0.5\mathcal{B}$ random ones and execute standard MPTS amortized evaluation and acquisition rule to obtain another $0.5\mathcal{B}$ ones from the rest random $\mathcal{B}$ identifiers, formulating the mixed $\mathcal{B}$ task batch for RL training. Such an operation makes RL over the MDP distribution stable in optimization.

### v5

@@ 5.3 Task Sampling Strategy Design
In robust active task sampling, existing strategies evaluate task batches to rank their difficulties in adaptation and then prioritize challenging subsets for optimization 39, 16, 17, 14, 40. Besides the expensive evaluation cost, these strategies are weak in the efficient exploration of the task space.
As Theorem 1 has established the theoretical foundation of approximately rank task difficulty, this necessitates the development of the risk learner from cumulated risk histories. With the model predictive results as amortized evaluation, specific rules can be incorporated into the acquisition function for active sampling. Meanwhile, it is fascinating for the risk learner to evaluate the machine learner’s adaptation to arbitrarily many tasks with minimal computational cost. Hence, we can easily enlarge the pseudo batch size $\hat{\mathcal{B}}$ for more selection candidates and exploit the epistemic uncertainty from the risk learner, encouraging more exploration in the task space.
@@
After obtaining $p_{\bm{\psi}}(\ell|\bm{\tau},H_{1:t})$, we draw up a batch sampling strategy on the basis of its quantified statistics. The criteria resembles the acquisition function in classical Bayesian optimization (BO), which includes a collection of available evaluation principles, such as expected improvement 52, output information theoretical index 53 or UCB 31.
However, it is also necessary to clarify that the search space is on the sequential task batch instead of machine learners’ parameters, which differs from the ultimate purpose in BO. Central to our approach is the principle of optimism in the face of uncertainty 54. We consider the difficult task’s prioritization for robustness and the epistemic uncertainty as pivotal elements in developing acquisition functions. The grounds behind this idea are that (i) the subset with the worst performance deserves extra attention in optimization for adaptation robustness, and (ii) task regions with high predictive uncertainty tend to be underexplored in the last few iterations.
As a result, we present the acquisition function built on the UCB principle 31:
@@
Then, the Simulate-Rank-Filter operation in Eq. (2)c arrives at the task batch for $(t+1)$-th iteration, i.e., $\mathcal{T}_{t+1}^{\mathcal{B}}=\arg\max_{\mathcal{T}^{\mathcal{B}}\subseteq\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}:|\mathcal{T}^{\mathcal{B}}|=\mathcal{B}}\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$. This characterizes the step of the active subset selection from $\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}$, the randomly sampled identifier candidate set with $|\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}|=\hat{\mathcal{B}}$. In an implementation, we still perform random sampling from $p(\tau)$ and forecast the task-wise acquisition score $a(\cdot)$ from the risk learner. Candidates in $-\mathcal{B}$ acquisition scores are screened to formulate the task batch $\mathcal{T}_{t+1}^{\mathcal{B}}$ for episodic optimization, as illustrated in Fig. 1d. These steps approximately solve Eq. (2)c and obtain $\mathcal{T}_{t+1}^{\mathcal{B}}$ in a heuristic way.
@@ 5.4 Sequentially Optimize the Adaptive Machine Learner
Given the screened $\mathcal{T}_{t+1}$, we execute optimization to update the machine learner’s parameters. The task-specific adaptation risk in $(t+1)$-th iteration is written as $\ell_{t+1,i}(\bm{\theta})$ for the selected task $\tau_{i}$. The developed MPTS is agnostic to any-shot learning methods, and the following includes the standard update rule for zero-shot, few-shot, and SFT scenarios.
@@
The zero-shot setup does not require the support dataset to identify the task. Hence, taking the vanilla DR 55 as an instantiation, we can obtain the update rule as:
@@
Still, we take the typical optimization-based method MAML 29 as an instantiation and provide the update rule as follows:
@@
Machine Learner Updates in SFT:
Here, we take finetuning pretrained models to downstream tasks 56 as an instantiation. In this case, each data point $[\bm{x},\bm{y}]$ can be viewed as a task with either its embedding $\bm{\tau}$ or $\bm{x}$ as the task identifier. Then the model update rule can be:
@@ 5.5 Overall Algorithm and Interpretation
Implementation Pipelines.
Here, we write the general form of MPTS in Algorithm 1, where the past risk episodes are reused to train risk learner and serve the active subset selection. We also provide some implementation examples by putting all the ingredients and optimization recipes together in the zero-shot, few-shot, and SFT scenarios. See Supplementary Notes in Algorithm 2-7 for details. Since the first iteration in Algorithm 2/4/6 does not involve active sampling, due to no latest history, and the task batch follows the standard random sampling setup.
@@
Output : Selected identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
@@
       $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
       $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
@@
// Simulating Zero-shot, Few-shot, Adaptation and SFT Results
@@
// Active Subset Selection from Predicted Results
@@
Return the screened identifier subset $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
@@
Intuitively, MPTS resembles model predictive control (MPC) 57 when treating task sampling under some criteria as an optimal planning problem. In this case, the episodic learning process specifies an underlying dynamical system for MPTS to predict with only one future time step in the simulation to assess the influence of selecting the task batch, and the feedback as exact adaptation risk information further helps improve the episodic risk prediction system. In addition, through the lens of sequential decision-making, we can interpret the optimization pipeline of MPTS from the actor-critic framework in RL 58. In detail, the risk learner works as the critic that predicts adaptation performance in the task $\tau$ given a fixed machine learner. Accordingly, the actor plays the role of selecting the task batch from the acquisition function and then executing the machine learner’s optimization. These two roles are entangled in the MPTS pipeline to achieve robust yet efficient adaptation.
@@
Supplementary Notes for Model Predictive Task Sampling
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
@@
// Active Subset Selection from Predicted Results
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
@@
// Active Subset Selection from Predicted Results
@@
Algorithm 6 MPTS for Pretrained Model Finetuning
@@
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (15b);
@@ Appendix A Quick Guideline to MPTS
Task episodic learning serves as a cornerstone in developing adaptive models by structuring diverse, context-rich learning experiences. One of the pivotal insights underpinning this process is the neural scaling law, which establishes a relationship between task volume, model complexity, and computational resources, offering a principled insight into training foundation models at a certain budget. Recent viewpoints have also shed light on the importance of task quality 59, 15, 60, 16, 18, 61, 62, 63, prompting innovative data curation strategies to refine datasets for pretraining, meta-training, and post-training. Evidence suggests that carefully curated data can significantly reduce task sampling complexity, decrease computational demands, and enhance robustness against distributional shifts—sometimes achieving these goals simultaneously. Despite these advancements, a practical operation such as Evaluate-Rank-Filter still faces challenges associated with costly evaluations from intensive task queries, computational overhead, and massive annotations. Addressing these bottlenecks remains essential to fully realize the potential of task episodic learning in robust efficient foundation model training.
Figure 8: Risk Predictive Module in MPTS for Active Subset Selection. MPTS adopts a predict-then-optimize strategy and uses a predictive module in green to obtain the preferred task subset. While the traditional method in blue exhausts $\hat{\mathcal{B}}$ tasks in construction and evaluation to filter preferred subset.
@@
The involvement of the risk learner inevitably brings extra computational overhead in optimization. However, the risk learner used in this work is lightweight with the model complexity $\mathcal{O}(|\bm{\phi}|+|\bm{\psi}|)<<\mathcal{O}(|\bm{\theta}|)$. We can roughly estimate these extra computations that arise from the predictive model as $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|)T_{\text{MPTS}}\big{)}$ throughout the training phase. Moreover, the computational and task evaluation complexities of different methods are estimated in Table 1. Compared with DRM, MPTS retains more computational and task efficiency when the filtering ratio $\hat{\alpha}$ is high, and the machine learner $\bm{\theta}$ is largely given similar convergence iteration steps.
@@
Among MPTS’s core components, the risk learner works to predict the adaptation risk values based on historical information and further serves the calculation of acquisition functions. Importantly, this work investigates the feasibility and effectiveness of risk predictive strategies and does not impose rigid constraints on the form of the risk learner $p(\ell|\bm{\tau},H_{1:t})$ too much in modeling. The design of this risk learner $p(\ell|\bm{\tau},H_{1:t})$ must meet several criteria: it is tractable in optimization, can process historical risk information, and offers uncertainty in prediction.
A series of candidate probabilistic models exist that probably apply to adaptation risk modeling. One alternative choice can be the Gaussian process 64, which provides an analytical form of the predictive distribution. However, its implementation (i) is less scalable in the case of relatively higher dimensional task identifiers, (ii) holds the cubic runtime complexity in obtaining the predictive covariance matrix, (iii) is sensitive to kernel selection, coupled with limited expressiveness of the Gaussian distribution in learned risk functions. Hence, for simplicity and computational efficiency, we adopt the basic VAE-like model and execute a handful of gradient updates to train the risk learner. We leave more advanced risk learner modeling for future exploration.
@@
This work relates to active sampling and Bayesian optimization. The purpose of BO 65 is to sequentially find a global optimum of a black-box function $f(\bm{x})$ expensive to evaluate in $\mathcal{S}$, namely $\bm{x}_{*}=\arg\max_{\bm{x}\in\mathcal{S}\subset\mathbb{R}^{d}}f(\bm{x})$.
In each iteration $t=1,\dots,T$, the BO method actively queries $\bm{x}_{t}$ to evaluate $f(\bm{x}_{t})$, yeilding an output $\ell_{t}=f(\bm{x}_{t})+\epsilon$ with a white noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Due to the high cost of function evaluation, the key to BO is constructing a surrogate model to guide the data point to query. The resulting acquisition function 66 works as an active sampling objective to maximize and obtain the candidate $\bm{x}_{t}$ based on the previous sequence. BO requires limited function evaluations as observations and exploits the correlations in queried data points. These properties make it more theoretically data efficient than random or grid search in seeking the optimal solution 67. This work differs from standard BO as task episodic learning is not the optimal parameter search problem.
@@
The main paper provides the workflow of MPTS in Algorithm 1. For separate scenarios, we attach detailed pseudo algorithms as follows. These illustrated Algorithms are in the context of supervised learning. Regarding RL scenarios, such as meta RL and DR, there is a slight modification for MPTS. As simply picking up worst-case MDPs restricts the task subspace in optimization 17, we adopt the mixture of the identifier subset from the random sampler and the identifier subset from the MPTS sampler. For example, in meta RL, with the pseudo batch size ^$\hat{\mathcal{B}}=1.5\mathcal{B}$, there $1.5\mathcal{B}$ identifier candidates from the random sampler. We retain $0.5\mathcal{B}$ random ones and execute standard MPTS amortized evaluation and acquisition rule to obtain another $0.5\mathcal{B}$ ones from the rest random $\mathcal{B}$ identifiers, formulating the mixed $\mathcal{B}$ task batch for RL training. Such an operation makes RL over the MDP distribution stable in optimization. See the open-source code for more RL details.